# Isotropic cosmic birefringence from string axion domain walls without cosmic strings, and DESI results

### v2

@@ 4 Summary
Considering the cosmological domain wall problem, domain walls of the string axion should be formed well after the recombination (Fig. 1). Such late formation can be realized if the axion has a hilltop initial condition (Fig. 2). As a result, the isotropic CB angle is robustly predicted as $\beta\approx 0.21c_{\gamma}$ deg, which agrees well with the recent ACT (and combined) result for $c_{\gamma}\sim 1$. While anisotropic CB is suppressed for the CMB photons from the recombination epoch due to the delayed domain wall formation, it is present for photons emitted at a later epoch. In addition, the domain walls themselves can induce the CMB anisotropies and gravitational waves.
Furthermore, we explored the implications of this scenario for time-varying dark energy, motivated by the recent DESI DR2 results [61], which further supports a deviation from a cosmological constant in addition to DESI DR1 [63]. We find that the equation-of-state parameter of dark energy can fit better the DESI results than the $\Lambda$CDM for low redshifts, $z\lesssim 1$.
In summary, our scenario highlights the intriguing possibility that late-time formation of domain walls associated with string axions could simultaneously explain isotropic cosmic birefringence and time-varying dark energy. To further assess the viability and testability of this scenario, future investigations on the CMB anisotropies, anisotropic CB, and gravitational waves will provide valuable insights, which we leave for future research.
@@ Acknowledgments
This work is supported by JSPS Core-to-Core Program (grant number: JPJSCCA20200002) (F.T.), JSPS KAKENHI Grant Numbers 20H01894 (F.T.), 20H05851 (F.T. and W.Y.), 22H01215 (W.Y.), 22K14029 (W.Y.), 23KJ0088 (K.M.), and 24K17039, (K.M.), Graduate Program on Physics for the Universe (J.L.) and JST SPRING Grant Number JPMJPS2114 (J.L.), Incentive Research Fund for Young Researchers from Tokyo Metropolitan University (W.Y.). This article is based upon work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology).

### v3

@@ 4 Summary
Considering the cosmological domain wall problem, domain walls of the string axion should be formed well after the recombination (Fig. 1). Such late formation can be realized if the axion has a hilltop initial condition (Fig. 2). As a result, the isotropic CB angle is robustly predicted as $\beta\approx 0.21c_{\gamma}$ deg, which agrees well with the recent ACT (and combined) result for $c_{\gamma}\sim 1$. While anisotropic CB is suppressed for the CMB photons from the recombination epoch due to the delayed domain wall formation, it is present for photons emitted at a later epoch, e.g., reionization. In addition, the domain walls themselves can induce the CMB anisotropies and gravitational waves.
[…]
In summary, our scenario highlights the intriguing possibility that late-time onset of string axion oscillations and the subsequent formation of domain walls could simultaneously explain isotropic CB and time-varying dark energy. To further assess the viability and testability of this scenario, future investigations on the CMB anisotropies, anisotropic CB, and gravitational waves will provide valuable insights, which we leave for future research.
@@ Acknowledgments
This work is supported by JSPS Core-to-Core Program (grant number: JPJSCCA20200002) (F.T.), JSPS KAKENHI Grant Numbers 20H01894 (F.T.), 20H05851 (F.T. and W.Y.), 22H01215 (W.Y.), 22K14029 (W.Y.), 23KJ0088 (K.M.), and 24K17039, (K.M.), Graduate Program on Physics for the Universe (J.L.) and JST SPRING Grant Number JPMJPS2114 (J.L.), Selective Research Fund for Young Researchers from Tokyo Metropolitan University (W.Y.). This article is based upon work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology).